OpenStudy (anonymous):

it's really hard to say mine question so if u understand please help me.... x squared +8x+20=0

OpenStudy (amistre64):

x^2 +8x +20 = 0 ; see that wasnt hard

OpenStudy (amistre64):

10 2 5 4 gonne be ugly with sqrts and all

OpenStudy (anonymous):

but i kinda want to know how you got the answer

OpenStudy (amistre64):

the answer, if there is one; might best be gotten by teh quadratic formula

OpenStudy (amistre64):

\[\frac{1}{2a}(-b\pm \sqrt{b^2-4ac})\]

OpenStudy (anonymous):

yes i know but then i end up with x=-8 plus or minus radical -16 over 2